### v3

# Hypo3D: Exploring Hypothetical Reasoning in 3D

Abstract
The rise of vision-language foundation models marks an advancement in bridging the gap between human and machine capabilities in 3D scene reasoning. Existing 3D reasoning benchmarks assume real-time scene accessibility, which is impractical due to the high cost of frequent scene updates. To this end, we introduce Hypothetical 3D Reasoning, namely Hypo3D, a benchmark designed to evaluate models’ ability to reason without access to real-time scene data. Models need to imagine the scene state based on a provided change description before reasoning. Hypo3D is formulated as a 3D Visual Question Answering (VQA) benchmark, comprising 7,727 context changes across 700 indoor scenes, resulting in 14,885 question-answer pairs. An anchor-based world frame is established for all scenes, ensuring consistent reference to a global frame for directional terms in context changes and QAs. Extensive experiments show that state-of-the-art foundation models struggle to reason effectively in hypothetically changed scenes. This reveals a substantial performance gap compared to humans, particularly in scenarios involving movement changes and directional reasoning. Even when the change is irrelevant to the question, models often incorrectly adjust their answers. The code and dataset are publicly available at: https://matchlab-imperial.github.io/Hypo3D.
Machine Learning, ICML
Figure 1: Overview of the Hypo3D benchmark. ① Examples of five context change types. ② Sample questions, including scale-based and direction-based questions requiring spatial reasoning, as well as semantic questions, all of which have open-ended answers. ③ The radar chart highlights a notable performance gap between models and humans, especially in direction-based questions.
1 Introduction
“Imagination is more important than knowledge.”
– ALBERT EINSTEIN
Artificial General Intelligence (AGI) aims to replicate the full spectrum of human cognitive abilities (goertzel2014artificial; rayhan2023artificial). Reasoning, a core cognitive ability, is a major research focus. The emergence of vision-language foundation models (hong20233d; liu2024visual; wang2024qwen2; fu2024scene; mao2024opendlign) has marked a significant step toward narrowing the gap between human and machine reasoning, enabling progress from text summarization (lewis2019bart; zhang2020pegasus) to complex scene understanding (huang2023embodied; driess2023palm).
These models are typically evaluated under the assumption that all knowledge about perception is immediately accessible during reasoning. For instance, in the evaluation of model performance in 3D scene reasoning, existing benchmarks (ye20213d; azuma2022scanqa; ma2022sqa3d; linghu2024multi; zhang2024spartun3d) presuppose real-time availability of scene data (e.g., point clouds). Yet, this assumption does not always hold, as real-world scenes are dynamic, and maintaining up-to-date 3D scenes is challenging. Unlike 2D image capture, 3D scene collection demands specialized equipment, extended scanning times, and a complex reconstruction process for accurate geometric representation (niessner2013real; daneshmand20183d). Thus, immediate perceptual knowledge for reasoning is unavailable in many real-world cases. Imagination, rooted in prior knowledge, allows humans to overcome such limitations by deducing missing details and approximating reality. Even without direct visual input, humans can mentally simulate changes in a scene and reason about the changed scene in their minds. This ability, known as mental imagery (pylyshyn2002mental; moulton2009imagining), is crucial to human intelligence and raises a critical question: Can current foundation models employ imagination to fill perceptual knowledge gaps and enhance reasoning?
In response to this question, we propose a new reasoning concept: hypothetical reasoning. It evaluates models’ ability to reason without immediate perceptual knowledge, requiring models to formulate reasonable hypotheses to bridge knowledge gaps. As an initial attempt, this study focuses specifically on hypothetical reasoning in 3D scenes, referred to as Hypo3D. As shown in Figures 1 and 2, the Hypo3D task follows a 3D visual question-answering format but cannot be solved using the given scene alone. Instead, models need to rely on an accessible context change description to imagine the current scene state after the change and adjust their answers accordingly. Based on the Hypo3D task, we constructed a dataset comprising 7,727 context changes and 14,885 question-answer pairs across 700 indoor scenes. As shown in Figure 1, these context changes span five categories: (1) Movement Change, involving geometric transformations like translation or rotation; (2) Removal Change, taking away objects; (3) Attribute Change, modifying object properties such as color and state; (4) Addition Change, introducing new objects; and (5) Replacement Change, substituting existing objects with new ones.
The questions in the dataset range from simple proximity and relative position queries to intricate path reasoning and navigation tasks, broadly categorized into three types: (1) scale-based, focusing on proximity and size relationships; (2) direction-based, requiring reasoning about directional terms; and (3) semantic, highlighting scene semantics with minimal spatial reasoning. In constructing direction-based questions, we observe that existing 3D reasoning datasets reveal ambiguities in defining directional terms in 3D. Early datasets (ye20213d; azuma2022scanqa) employ object-centric definitions, which lead to confusion when dealing with symmetrical or amorphous objects, such as round tables or cushions. Recent datasets (ma2022sqa3d; linghu2024multi; zhang2024spartun3d) define directions relative to the observer. This strategy struggles to describe global directional relationships and introduces inconsistencies when multiple observers are involved. Towards this, Hypo3D establishes a world frame anchored to reference objects in each scene, ensuring that all directional terms are defined relative to a consistent global reference frame.
Extensive experiments on ten foundation models reveal a substantial performance gap between models and humans on the Hypo3D task, particularly in movement changes and directional reasoning. Surprisingly, closed-source models (e.g., GPT-4o (openai2024gpt4o)) do not outperform open-source alternatives. Furthermore, the models frequently hallucinate context changes, modifying their answers even when those changes are irrelevant to the posed question. Notably, their performance consistently degrades when required to imagine a scene after changes before reasoning, as compared to directly reasoning with the provided scene. These findings highlight a key limitation of current models: the inability to leverage imagination to infer missing perceptual knowledge and reason hypothetically.
Figure 2: Example of hypothetical reasoning in a 3D scene. Given a 3D scene and an anchor-based frame description (Scene Orientation), models first align the scene to the specified frame. Then, based on a context change description and a question, models hypothetically modify the aligned scene and answer questions about the changed scene. Various models (e.g., LLMs, 2D VLMs, 3D VLMs) can tackle this task using corresponding scene representations, including scene captions, top-view maps, point clouds, and egocentric RGB-D videos.
2 Related Work
2.1 3D Scene Update
The evolving nature of 3D environments requires continuous scene updates for effective understanding, particularly in autonomous driving (yurtsever2020survey) and robotic navigation (wong2000scene). A fundamental approach to scene updating involves reconstructing the entire scene from scratch. Advances in 3D reconstruction, such as multi-view stereo (seitz2006comparison; jing2025match), depth sensor-based (zollhofer2014real), and volumetric methods (newcombe2011kinectfusion), have improved reconstruction fidelity. But current methods struggle to balance accuracy, efficiency, and scalability (niessner2013real), making reconstruction primarily suitable for situations involving major changes. Recently, radiance-based 3D scene editing approaches, such as NeRF and 3D Gaussian Splatting (3DGS), have emerged as more efficient alternatives for incremental scene updates. NeRF-based methods (liu2021editing; kania2022conerf) were limited to basic object edits and struggled with complex, cluttered scenes (ye2025gaussian). 3DGS-based methods enable finer control over scene content, including geometry (huang2024sc; waczynska2024games; ye2025gaussian), texture (chen2024gaussianeditor), and lighting (gao2025relightable). However, they struggle to disentangle these components and require costly re-optimization, limiting editability and efficiency (wu2024recent).
Therefore, accurate and efficient 3D scene updates remain a challenge, making real-time scene acquisition not always feasible. This highlights the need for hypothetical reasoning.
2.2 3D Visual Question Answering
3D Visual Question Answering (3D VQA), a key task for evaluating 3D reasoning, has advanced with the rise of Vision-Language Models (VLMs) (hong20233d; liu2024visual; anthroptic2024claude3.5sonnet). By integrating vision encoders with Large Language Models (LLMs) (peng2023instruction; bai2023qwen), VLMs enable multimodal perception of 3D scenes using inputs like top-view maps and point clouds. As models advance, there is growing interest in developing more comprehensive benchmarks to better assess their 3D reasoning capabilities. Initial 3DQA dataset (ye20213d), derived from ScanNet (dai2017scannet), introduced 6K manually annotated QA pairs for scene-level reasoning. ScanQA (azuma2022scanqa) expanded this with 41K QA pairs using automated question generation and human refinement. Qian et al. (qian2024nuscenes) further extended 3D VQA to autonomous driving scenarios, offering domain-specific challenges. SQA3D (ma2022sqa3d) introduced “situated reasoning,” requiring models to contextualize answers based on an agent’s position and orientation. MRSA (linghu2024multi) and SPARTUN3D (zhang2024spartun3d) further scaled SQA3D, adding multimodal inputs such as images for richer situational context.
Unlike previous benchmarks, where answers are derived fully from the given scene, Hypo3D hypothetically applies a change to the scene and derives answers from the changed scene, increasing the hallucination risk.
3 Hypo3D Benchmark
Figure 3: Dataset Generation Pipeline. The Hypo3D collection pipeline consists of five stages: Stage ① curates scenes (50 hours per person), Stage ② defines world frames (10 h/p), Stages ③ and ④ collect context changes and QA descriptions from human annotators (thousands of hours) and LLM, and Stage ⑤ conducts grammar checks and filters data based on semantic similarity. (25 h/p).
This section first defines the Hypo3D task and explains how humans, assisted by LLMs, generate high-quality context changes and QAs. It then presents the dataset statistics.
3.1 Task Definition
A task instance in Hypo3D is formulated as a tuple ⟨$S$, $F$, $c$, $q$⟩, where $S$ denotes the scene representation; $F$ specifies the world frame defined by anchor objects in $S$ that standardizes all directional terms in the task; c describes a context change to be applied to the scene; and q denotes a question. The task is to first rotate $S$ into $\tilde{S}$ so that the anchor object is located in the orientation described in $F$, followed by computing an answer $a$ to $q$ after applying $c$ to $\tilde{S}$. Figure 2 provides an example of hypothetical reasoning in a 3D scene.
3.2 Dataset Generation Pipeline
As depicted in Figure 3, the Hypo3D benchmark was constructed through a rigorously designed five-stage pipeline. The initial Stage ① involves a comprehensive 50-hour manual curation process to collect and standardize diverse 3D scene representations from established datasets. This is followed by Stage ②, a meticulous 10-hour manual annotation phase where each scene is assigned an anchor-based world frame to ensure spatial consistency. The core data generation occurs in Stages ③ and ④, where we implemented a hybrid approach combining human expertise with LLMs to generate context changes and question-answer pairs. These critical stages demanded substantial human oversight, with over 1,000 hours dedicated to human refinement to guarantee exceptional data quality and contextual coherence. Finally, the pipeline concludes with Stage ⑤, a 25-hour post-processing phase that employs quality control measures to eliminate redundancies while maintaining grammatical precision.
Scene Curation. Each 3D scan $S$ in the Hypo3D benchmark can be represented through multiple modalities to facilitate reasoning, including point cloud, RGB-D video, top-view map, and scene caption. The point clouds and RGB-D videos are directly curated from the original scene datasets. Two variants of top-view maps are generated: non-semantic and semantic. Non-semantic maps are generated by positioning a simulated orthographic camera above the scene to produce photorealistic top-down renderings. Semantic maps augment non-semantic ones by incorporating hoverable semantic labels, with each label precisely positioned at the centroid of the bounding box corresponding to its object. An example semantic map can be seen in Figure 2. Unlike their non-semantic counterparts, semantic maps reduce object recognition errors in models, prioritizing the evaluation of hypothetical reasoning performance. Scene captions are textual descriptions detailing the attributes of objects and their relationships within the scenes.
Anchor-Based World Frame Annotation. The same scene in 3D can be captured from different viewpoints in any orientation. Thus, it does not have a clearly defined, fixed reference frame. Here, we defined the world coordinate frame textually based on the locations of anchor objects in each scene, such as: “The desk is located to the left of the scene”. When selecting anchor object candidates, visual prominence is prioritized over the extent to which they cover multiple sides of the scene. This is because, conventionally, defining a single primary orientation (e.g., “front”) is sufficient to infer the remaining orientations (“left”, “right”, “back”) using the right-hand rule (hamilton2008lectures), ensuring a consistent and unambiguous spatial frame.
Context Change Preparation. The Hypo3D benchmark defines five distinct types of object-level context changes, illustrated in Figure 1. For a scene $S$ with an object list $O=\{o_{1},o_{2},\dots,o_{n}\}$, each change type is defined as follows:
  1. (1)
    Movement changes relocate objects $O_{m}\subseteq O$ to new positions or orientations within the scene $S$.
  2. (2)
    Removal changes eliminate objects $O_{r}\subseteq O$ from $S$, resulting in an updated object list $\hat{O}=O\setminus O_{r}$.
  3. (3)
    Attribute changes modify properties of a subset of objects $O_{att}\subseteq O$, such as color, material, or state.
  4. (4)
    Addition changes introduce new objects $O_{a}$ into $S$, updating the object list to $\hat{O}=O\cup O_{a}$.
  5. (5)
    Replacement changes replace objects $O_{rp}\subseteq O$ with new objects $O_{a}$, resulting in an updated object list $\hat{O}=(O\setminus O_{rp})\cup O_{a}$.
As illustrated in Figure 3 ③, human annotators and GPT-4o initially generate raw context changes uniformly to ensure data diversity. Human data is collected via crowdsourcing on the CloudResearch platform. [^1] CloudResearch, Connect Platform, https://www.cloudresearch.com. For each scene aligned to the world frame $\tilde{S}$ with an associated object list $O$, annotators create a set of distinct descriptions $C$ for a specified change category, following the rules below:
  1. (1)
    Each object $o_{i}\in O$ referenced in change $c_{i}\in C$ must have a uniquely specified location if it appears multiple times in the scene $\tilde{S}$.
  2. (2)
    Each context change $c_{i}$ must be spatially feasible within the layout of $\tilde{S}$.
  3. (3)
    Each $c_{i}$ must be independent, meaning it is derived solely from the original scene $\tilde{S}$ and does not rely on any version of $\tilde{S}$ modified by $c_{j}$, where $i\neq j$.
For GPT-4o, context changes are generated using the semantic top-view map of $\tilde{S}$ and a textual prompt that follows the same criteria as those provided to human annotators. The detailed prompts and human guidelines for raw change generation are provided in Appendix LABEL:app:A.
Finally, each raw change $c$ generated by humans and GPT-4o is edited by an independent group of human reviewers, producing a refined version $c^{\prime}$.
Figure 4: Dataset Statistics. ① Word cloud representing context change descriptions. ② Frequency distribution of context change types across 7,727 instances. ③ Distribution of question types across change categories, with question frequency consistently highest for scale-based, then direction-based, and finally semantic.
Question-Answer Preparation. Raw questions are initially generated by GPT-4o, as illustrated in Figure 3 ④. Seven prompt templates, each corresponding to a specific question type, are designed for each type of context change. Each template includes the object list $O$ in the scene $S$, the context change $c$, example questions, and a question type description (e.g., ask for the position of the changed object relative to other objects in the scene). In total, eleven unique question types are evenly distributed across the context change types (see Appendix LABEL:app:A.2). For each context change $c$, 21 raw questions are generated, denoted as $Q_{c}$.
Human reviewers then refine $Q_{c}$ to produce a new question set $Q_{c}^{\prime}$. During this process, only a small portion of raw questions are retained. Specifically, 91% of the questions are filtered out, and the remaining questions undergo additional editing to strictly ensure $Q_{c}^{\prime}$ satisfies the following criteria: (1) each $q\in Q_{c}^{\prime}$ can only be answered by combining $S$ and $c$, as neither is sufficient on its own; (2) the answer to each $q$ must be potentially impacted by the change $c$; (3) each $q$ has a unique and unambiguous answer; and (4) answers cannot be inferred from commonsense knowledge (e.g., a bed is larger than a pillow). Finally, all questions in $Q_{c}^{\prime}$ are annotated with concise human-provided answers $A_{c}^{\prime}$ and reclassified into general types to evaluate models based on their reasoning capabilities: scale-based, direction-based, and semantic questions, as shown in Figure 1. Scale-based questions assess spatial reasoning related to proximity and size perception, direction-based questions focus on orientation understanding, and semantic questions evaluate the model’s ability to interpret object attributes with minimal spatial reasoning. Notably, questions can belong to multiple types if they require diverse reasoning.
Post-processing. To ensure data diversity, SBERT (reimers2019sentence) is used to remove semantically similar descriptions. It encodes the refined context changes $C^{\prime}$ and questions $Q^{\prime}$, producing embeddings $E_{C^{\prime}}$ and $E_{Q^{\prime}}$, respectively. Context changes and questions are filtered by excluding pairs with cosine similarity $\text{Sim}(e_{i},e_{j})>0.8$ for $e_{i},e_{j}\in E_{C^{\prime}}$, and $\text{Sim}(e_{k},e_{v})>0.8$ for $e_{k},e_{v}\in E_{Q^{\prime}}$. The remaining context changes, questions, and corresponding answers are further refined for grammatical accuracy using GPT4-Turbo, resulting in descriptions for the final dataset $\tilde{C}$, $\tilde{Q}$, and $\tilde{A}$.
3.3 Statistics
The Hypo3D benchmark comprises 700 unique scenes, with 500 sourced from the ScanNet (dai2017scannet) dataset and 200 from the 3RScan (wald2019rio) dataset, randomly sampled from their respective sources. The dataset includes 7,727 context changes and 14,885 question-answer pairs. On average, a context change description contains 13.62 words, a question description contains 13.69 words, and an answer contains 1.28 words. The word cloud in Figure 4 ① highlights that the most frequent words in context change descriptions are verbs representing change actions, such as “moved,” “removed,” and “replaced”. Figure 4 ② illustrates the distribution of context change types, with movement changes being the most frequent, as they often result in more pronounced scene layout rearrangements. The bar charts in Figure 4 ③ show that scale-based and direction-based questions, which require more spatial reasoning, constitute a larger portion of the dataset compared to semantic questions.
4 Experiments
4.1 Evaluation Protocol
Exact Match (EM) and Partial Match (PM) are the metrics used for evaluation. EM measures the percentage of model-predicted answers that exactly match the ground-truth answers. PM quantifies the percentage of overlapping words between the predicted answers and the ground truth. For the same model, PM is typically higher than EM, as it accounts for partial correctness. When computing EM and PM, both the predicted and ground-truth answers are normalized by lowercasing, removing punctuation, and aligning semantically similar terms (e.g., “left” and “west”).
Table 1: EM and PM accuracy of ten foundation models and human evaluators on Hypo3D. The highest model performance for each type of context change is in bold, while the best-performing model within each family is underlined.